11. Let A= S- Find a matrix P such that P-IAP is a diagonal matrix.
Answer #1

A 179] (A-2)=0 and y 1-1-) (2-476-1-2) - 30 -2+2-27 +2_4-0 arx-620 2-3x+22-620 (1-3)+2(2-3) (4+2) (2-3) 2.0 = 3,-2 133 4 4 1